Test file for hyperref and color.

Contents

1 HyperTEX tests 2 1.1 A subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Yet another subsection . . . . . . . . . . . . . . . . . . . . . . 2

2 Color tests 3

3 Background tests 4

1 HyperTEX tests

1.1 A subsection

See color section 2.

1.2 Yet another subsection

See background section on page 4.

Here’s some nasty links and targets.

This link has a target in the middle. Note that an extra space occurs after the target; this can be removed by using hyperref’s “nesting” option.

This target has a link in the middle.

Links of different sizes: [tiny] [huge]

Here’s a link that spans a line break: This is a very very very very very very very very very very long link. And here’s an even longer link: This is a very very very very very very very very very very very very very very very very

very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very very long link.

And here’s a link that spans a page break: This link continues...

... this link starts on the previous page, so is ignored.

2 Color tests

Note that colored text is not dithered (nor are colored links).

This sentence should be black. This should be red. This should be green.

Black. Blue. Cyan. Magenta. Yellow. Back to Magenta. Back to Cyan. Back to Blue. Back to black.

Light gray. Dark gray. Purple. No idea what this color is! Nor this.

Black text on a red background.

Black text on a yellow background with red frame.

Now let’s test the use of color names defined in color.pro: GreenYellow Yellow Goldenrod Dandelion Apricot Peach Melon YellowOrange Orange BurntOrange Bittersweet RedOrange Mahogany Maroon BrickRed Red OrangeRed RubineRed WildStrawberry Salmon CarnationPink Magenta VioletRed Rhodamine Mulberry RedViolet Fuchsia Lavender Thistle Orchid DarkOrchid Purple Plum Violet RoyalPurple BlueViolet Periwinkle CadetBlue CornflowerBlue MidnightBlue NavyBlue RoyalBlue Blue Cerulean Cyan ProcessBlue SkyBlue Turquoise TealBlue Aquamarine BlueGreen Emerald JungleGreen SeaGreen Green ForestGreen PineGreen LimeGreen YellowGreen SpringGreen OliveGreen RawSienna Sepia Brown Tan Gray Black White

The text on the rest of the page is blue (but not the page number).

3 Background tests

The background for this page should be yellow. Text should still be blue. Back to black. This text should be white.

The background should still be yellow. This page is deliberately... ...off the paper.

The background is now black and the text is white. 6